Calculate the degree of unsaturation for the compound c24h26no3cl3
1
Expert's answer
2019-04-04T07:40:45-0400



Task:

Calculate the degree of unsaturation for the compound C24H26NO3Cl3.


Solution:

The degree of unsaturation (DU) is a calculation that determines the total number of rings and π bonds.

rings + π bonds = DU = N(C) - N(H)/2 + N(elements of V group)/2 - N(elements of VII group)/2 + 1

C24H26NO3Cl3

N(C) = 24

N(H) = 26

N(elements of V group) = N(N) = 1

N(elements of VII group) = N(Cl) = 3

Then,

DU =  N(C) - N(H)/2 + N(Cl) - N(N) = 24 - (26 / 2) + (1 / 2) - (3 / 2) + 1 = 24 - 13 + 0.5 - 1.5 + 1 = 11

The degree of unsaturation (DU) = 11